# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v2

@@ 3.2 Method: CLIMB-3D
Overview. As illustrated in Fig. 2 and introduced in Sec. 3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy. In each phase, the model is exposed to a carefully selected subset of the dataset, designed to simulate real-world scenarios discussed in Sec. 3.3. Naïvely training on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets previously acquired knowledge when learning new tasks. To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], inspired by 2D [rebuffi2017icarl, li2017learning, aljundi2018memory] and 3D [boudjoghra20243d] incremental setting, to mitigate forgetting by storing a small set of representative samples from earlier phases. However, ER alone is insufficient for achieving robust performance. We therefore introduce two additional components: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases. Each of these components is described in detail below  [^2] For clarity, modifications introduced by each component – ER, PLG, and CBR – are highlighted in green, blue, and red, respectively..
@@
Exemplar Replay (ER). Inspired by Buzzega et al. [buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a small portion of data from earlier phases. At phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$ but also on a set of exemplars $\mathcal{E}^{1:t-1}$ collected from previous phases. These exemplars are accumulated incrementally, where $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1},$ and $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, with $|\mathcal{E}^{t}|$ as the exemplar replay size. We define the combined replay dataset as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training proceeds in two stages: first on the current data $\mathcal{D}^{t}$, followed by replay using $\mathcal{E}^{1:t-1}$. The updated objective function for CI-3DIS, based on LABEL:eq:ci3dis-loss, becomes:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|\mathbf{Y}^{t}|}\sum_{(m_{j}^{t},c_{j}^{t})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j}^{t},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j}^{t},\hat{c}_{j}^{t})\right).$ (3)
@@
While previous CI-3DIS approaches use large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints. Consequently, although exemplar replay (ER) aids knowledge retention, it alone proves insufficient for robust performance with a limited exemplar budget. We therefore propose PLG and CBM to better preserve past representations and enhance generalisation across phases.
@@
Pseudo-Label Generator (PLG). In the CI-3DIS setting, at phase $t$ (for $t>1$), although we no longer have access to the ground-truth labels for previously seen classes, the model from the previous phase, $\Phi^{t-1}$, remains available and retains knowledge acquired from past tasks. We use this model to generate pseudo-labels for previous classes, thereby providing approximate past supervision during training. This allows the current model $\Phi^{t}$ to retain past knowledge and reduces forgetting, even without access to original annotations.
@@
Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$ from $\mathcal{D}^{t}$, the previous model model $\Phi^{t-1}$ generates pseudo-labels for the previously learned classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity. However, we observe that using all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance. To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$. At the end, Using this supervision set, the model is optimized via the following updated objective:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}})\in{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}}$
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (4)
where the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from previous phases, enabling the model to retain prior knowledge while adapting to new classes.
@@
Class-Balanced Re-weighting (CBR). While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased toward frequently occurring classes, leading to forgetting of rare categories. This issue is further reinforced by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data. To alleviate this issue, we introduce a Class-Balanced Re-weighting (CBR) scheme that adjusts for class imbalance based on frequency of object occurrence. However, at task $t$, we only have access to the dataset $\mathcal{D}^{t}$ and its corresponding label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$. The datasets and class distributions from previous tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible – making it challenging to directly account for class imbalance across all seen tasks. Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
@@
During each training iteration, we apply the previous model $\Phi^{t-1}$ to the current input point cloud $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$ , from which we accumulate class-wise frequency statistics corresponding to prior tasks. Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$. At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the observed ground-truth label distribution $p^{t}(c)$ from the current task. The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$ weights for balanced pseudo-label selection and to tune current model to favour rare classes.
@@
During the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$. To promote balanced pseudo-label selection, we apply class-wise re-weighting using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability. The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ We then perform top-$K$ selection over both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively. The final supervision label set $\overline{\mathbf{Y}}^{t}$ is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels: ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$
Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$. To resolve this, we extend the class-balanced weights $w_{c}$ to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$, and denote the updated weights as $w_{c}^{\prime}$, which are then integrated into the training objective. After incorporating all components, the modified objective in Eq. 1 becomes the final updated objective used in CLIMB-3D:
@@
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}$
@@
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (5)
where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a loss-balancing hyperparameter. To address the class-imbalance across tasks, our CBR module jointly re-weights pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones. This unified strategy reduces bias and enhances performance, particularly for under-represented classes.

### v3

@@ 3.2 Method: CLIMB-3D
Overview. As shown in Fig.˜2 and formulated in Sec.˜3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy. In each phase, the model is exposed to a carefully curated subset of the dataset, simulating real-world scenarios discussed in Sec.˜3.3. Training naïvely on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets prior knowledge when learning new tasks. To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], a strategy inspired by 2D incremental methods [rebuffi2017icarl, li2017learning, aljundi2018memory] and recently extended to 3D settings [boudjoghra20243d], to mitigate forgetting by storing a small set of representative samples from earlier phases. However, ER alone is insufficient to achieve robust performance. To address this, two additional components are introduced: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases. Each of these components is described in detail below  [^1] For clarity, modifications introduced by each component (ER, PLG, and CBR) are highlighted in green, blue, and red, respectively..
@@
Exemplar Replay (ER). Inspired by Buzzega et al. [buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a subset of data from earlier phases. During phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$, but also on a set of exemplars $\mathcal{E}^{1:t-1}$, collected from previous phases. These exemplars are accumulated incrementally: $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1}.$ Here, $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, and $|\mathcal{E}^{t}|$ denotes the exemplar replay size. The combined replay dataset is denoted as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training consists of two stages: first, the model is trained on the current data, $\mathcal{D}^{t}$; then it is trained using exemplars from $\mathcal{E}^{1:t-1}$.
@@
While previous CI-3DIS approaches rely on large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints. Consequently, although ER aids knowledge retention, it is insufficient on its own for robust performance with a limited exemplar budget. Therefore, PLG and CBR are introduced to better preserve past representations and enhance generalisation across phases.
@@
Pseudo-Label Generator (PLG). In the CI-3DIS setting, during phase $t$ (where $t>1$), although ground-truth labels for previously seen classes are unavailable, the model from the previous phase $\Phi^{t-1}$, which preserves knowledge of past tasks, is retained. We use this model to generate pseudo-labels for previously seen classes, thereby providing approximate supervision during training. This allows the current model $\Phi^{t}$, to retain prior knowledge and reduce forgetting, even without access to ground-truth annotations.
@@
Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$, from $\mathcal{D}^{t}$, the previous model $\Phi^{t-1}$, generates pseudo-labels for the previously seen classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity. However, we observe that utilising all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance. To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$. Following this, the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from earlier phases, enabling the model to retain prior knowledge while adapting to new classes.
@@
Class-Balanced Re-weighting (CBR). While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased towards frequent classes, resulting in the forgetting of rare categories. This issue is further amplified by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data. To address this, we propose a Class-Balanced Re-weighting (CBR) scheme that compensates for class imbalance using object frequency statistics. At task $t$, only the current dataset $\mathcal{D}^{t}$ and its label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$ are available. Datasets and class distributions from earlier tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible, which makes it challenging to directly account for class imbalance across all previously seen tasks. Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
@@
During each training iteration, the previous model $\Phi^{t-1}$ is applied to the current input $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$, from which we accumulate class-wise frequency statistics for previously learned classes. Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$. At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the ground-truth label distribution $p^{t}(c)$ from the current task. The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise weights $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$, which guide balanced pseudo-label selection and help the model focus on rare classes.
@@
In the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ (as described in PLG) , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$. To promote balanced pseudo-label selection, class-wise re-weighting is applied using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability. The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ Top-$K$ selection is then applied to both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively. The final augmented target set, $\overline{\mathbf{Y}}^{t}$, is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels:
@@
${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$ (2)
Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$. To resolve this, the class-balanced weights $w_{c}$ are extended to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$. The updated weights, denoted as $w_{c}^{\prime}$, are used to re-weights both pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones. This unified strategy reduces bias and enhances performance across phases. The pseudo-code for constructing the supervision set in Eq.˜2, which incorporates both PLG and CBR components, is provided in Appendix A.
Final Objective. The final training objective of CLIMB-3D, incorporating ER, PLG, and CBR, is formulated in Eq.˜1 as:
@@
$\mathcal{L}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (3)
where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a hyperparameter for balancing the loss terms.